# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ Appendix A Appendix: PXP Model
$\hat{H}=\hat{S}_{0}\bigoplus\hat{T}+\hat{V},$ (S1)
where $\hat{S}_{0}$ and $\hat{T}$ denote the Hamiltonian of scar and thermal subsystems, respectively, while $\hat{V}$ represents the coupling between them. This general Hamiltonian includes a large class of QMBS models [11]. For concreteness, here we will focus on the spin Creutz-ladder (SCL) model introduced in the main text, where $\hat{S}_{0}$ takes the form of a free paramagnet Hamiltonian, to be specified below.
The effect of the thermal subspace $\hat{T}$ can be incorporated by defining the effective Hamiltonian for the scar subspace as $\hat{S}=\hat{S}_{0}+\hat{\Sigma}$, where $\hat{\Sigma}$ is the self-energy from the thermal subspace. If we neglect the coupling between the bath and the thermal subsystem, the resulting Liouvillean can be written as
@@
Here, the Liouvillean of the exact scar subspace is denoted by $\mathcal{S}_{0}$, while the self-energy and dissipator are
@@
The self-energy can be obtained via the non-equilibrium Green’s function as shown below. When $\mathcal{E}$ is zero, the problem reduces to the exact scar case. The effective dephasing strength, $\gamma_{\mathrm{eff}}$, is model-dependent. For example, in the SCL model, two neighboring sites in the real space are used to define a ‘spin’ in the scar subspace (i.e., $N=L/2$), thus we have $\gamma_{\mathrm{eff}}=2\gamma$.
In the SCL model, the QMBS Hamiltonian and its eigenstates are given by the Dicke states of a free collective spin of magnitude $N/2$:
$\hat{S}_{0}=J\sum_{j}\hat{\sigma}^{x}_{j},\quad\quad\quad\left|E_{n_{S}}\right\rangle=\binom{N}{n_{S}}^{-\frac{1}{2}}\sum_{i_{1}\neq\cdots\neq i_{n_{S}}}\bigotimes_{j=1}^{N}\left\{\begin{array}[]{ll}|+\rangle&j\in\left\{i_{1},\ldots,i_{n_{S}}\right\}\\ |-\rangle&\text{ otherwise }\end{array},\right.$ (S4)
where |±$|\pm\rangle_{j}=\big{(}|0\rangle_{j}\pm|1\rangle_{j}\big{)}/{\sqrt{2}}$ are the eigenvectors of a local $\hat{\sigma}_{j}^{x}$, and $n_{S}=0,1,\ldots,N$ is the number of |+ket|$\rangle$| + ⟩ appearing in the tensor product. We have written the eigenstates belonging to the maximal spin representation. The dephasing-free Liouvillean of the exact scar subspace is given by
$\mathcal{S}_{0}={-iJ\Big{(}\sum_{j}\hat{\sigma}^{x}_{j}\otimes\hat{I}-\hat{I}\otimes\sum_{j}\hat{\sigma}^{x}_{j}\Big{)}},$ (S5)
@@
$|\lambda_{(l,s)}^{(0)}\rangle\!\rangle=|E_{l+s}\rangle\otimes|E_{s}\rangle,\quad\quad\quad\quad\lambda_{l}^{(0)}=i(E_{l+s}-E_{s})=2ilJ.$ (S6)
The eigenmodes are labeled using indices $l=-N,-N+1,\cdots,N$ and $s=0,1,2,\cdots,N-|l|$. The layer indices can be split into two groups, i.e. $\{l\}=\{l^{\prime},l^{\prime\prime}\}$, where QMBS layers $l^{\prime\prime}=-N,-N+2,\cdots,N$ represent layers with spontaneous LPTS breaking process and $l^{\prime}=-N+1,-N+3,\cdots,N-1$ are the remaining layers.
A few comments are in order. The layers $l^{\prime\prime}=\pm N$ are special in that they contain a single level $s=0$. Consequently, as we show in Sec. A2 below, these QMBS Liouvillean eigenvalues remain on the imaginary axis for arbitrary values of the dephasing, while the other $l^{\prime\prime}$ sectors generally undergo LPTS breaking. Second, we note that our labeling of QMBS layers is unrelated with inversion symmetry quantum number. The latter is given by $p=(-1)^{l}$, therefore the LPTS-breaking states are found in different inversion-symmetric subspaces depending on the parity of $N$.
S1.1 A1. Self-energy due to the thermal subspace
The Hamiltonian in Eq. (S1) is generally non-integrable and impossible to solve exactly. Hence we employ the non-equilibrium Green’s function approach to model the effect of thermal bulk on the scar subspace, see Ref. [32] for details. This will allow us to obtain analytical insights into the interplay of scarring with the dephasing dissipation. The Green’s function of the scar subspace is expressed as
@@
where $\hat{\Sigma}(E)=\sum_{p_{\mathrm{S}}}\hat{\sigma}_{p_{\mathrm{S}}}(E)$ represents the self-energy matrix, with $\hat{\sigma}_{p_{\mathrm{S}}}(E)=\hat{\Gamma}_{p_{\mathrm{S}}}^{\dagger}\hat{G}_{0}(E)\hat{\Gamma}_{p_{\mathrm{S}}}\text{ and }\hat{\Gamma}_{p_{\mathrm{S}}}=\sqrt{\gamma_{p_{\mathrm{S}}}}\left|p_{\mathrm{S}}\right\rangle\left\langle p_{\mathrm{S}}\right|$. Here, $|p_{\mathrm{S}}\rangle$ denote the product states in the scar subspace, and $\gamma_{p_{\mathrm{S}}}$ represents the interaction strength between the $|p_{\mathrm{S}}\rangle$ and the thermal subspace. For the QMBS model we consider here, Eq. (S1), $\hat{S}_{0}$ corresponds to a hypercube and $|p_{\mathrm{S}}\rangle$ represent some of the computational basis states, while the remaining basis states $|p_{\mathrm{T}}\rangle$ span the thermal subspace. For example, in the SCL model the vertices of the hypercube are product states of dimers, $\{|p_{\mathrm{S}}\rangle\}=\{\otimes_{j=1}^{N=L/2}|d_{j}\rangle\}$, where each $|d_{j}\rangle$ can be either $|01\rangle_{j}$ or $|10\rangle_{j}$. The thermal states $|p_{\mathrm{T}}\rangle$ then involve the remaining basis states that contain $|00\rangle_{j}$ and $|11\rangle_{j}$ states of dimers.
@@
For the SCL model, the interaction strength $\gamma_{p_{\mathrm{S}}}$ is proportional to the hopping matrix element $h_{p_{\mathrm{S}}}=\sum_{p_{\mathrm{T}}}\left\langle p_{\mathrm{T}}\right|\hat{V}\left|p_{\mathrm{S}}\right\rangle$ between the scar $|p_{\mathrm{S}}\rangle$ and the thermal subspace $|p_{\mathrm{T}}\rangle$ [32]. The expression for $\hat{\Gamma}$ can be derived analytically:
$\hat{\Gamma}_{\mathrm{SCL}}=\sum_{p_{\mathrm{S}}}\sqrt{\gamma_{p_{\mathrm{S}}}}|p_{\mathrm{S}}\rangle\langle p_{\mathrm{S}}|=\sum_{p_{\mathrm{S}}}|p_{\mathrm{S}}\rangle\langle p_{\mathrm{S}}|\sqrt{\sum_{p_{\mathrm{T}}}\left\langle p_{\mathrm{T}}\right|\hat{V}\left|p_{\mathrm{S}}\right\rangle}=\sum_{p_{\mathrm{S}}}|p_{\mathrm{S}}\rangle\langle p_{\mathrm{S}}|\sqrt{\sum_{j=1}^{N-1}J_{\mathrm{x},j}\hat{P}^{j,j+1}_{p_{\mathrm{S}}}},$ (S9)
@@
where $\hat{P}_{p_{\mathrm{S}}}^{j,j+1}=\langle d_{j}|d_{j+1}\rangle$ represents the overlap between two neighboring dimers in $|p_{\mathrm{S}}\rangle$. In the following, we use this expression for $\hat{\Gamma}$ to evaluate the self-energy from the Dyson equation.
S1.2 A2. Degenerate perturbation theory for the Liouvillean
To explain the structure of Liouvillean QMBS spectra, we set up a degenerate perturbation theory framework. We consider the Liouvillean of the exact scar subspace, Eqs. (S5) and (S6), with eigenvalues
$\lambda_{l}^{(0)}=i(E_{n_{\mathrm{S}}}-E_{m_{\mathrm{S}}})=2ilJ,$ (S10)
where $l=n_{\mathrm{S}}-m_{\mathrm{S}}=-N,-N+1,\cdots,N-1,N$. The eigenvalues $\lambda_{l}$ are equally spaced along the imaginary axis, with the real part equal to zero.
In the regime $\|\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}+\mathcal{E}\|\ll\|\mathcal{S}_{0}\|$, we can use the Liouvillean perturbation theory, in which the traceless Liouvillean eigenvalues for the $l$th layer are approximately given by $\lambda^{\prime}_{l}\approx\lambda^{(0)}_{l}+\lambda^{(1)}_{l}$. Here, the zeroth-order eigenvalues are given by the eigenvalues of the unperturbed system in Eq. (S10), while the first-order corrections to the eigenvalues, $\lambda_{l}^{(1)}$, are obtained by diagonalizing the projection of $\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}+\mathcal{E}$ to the degenerate subspace of the $l$th layer, labeled by states $s=0,1,2,\cdots,N-|l|$.
First, we consider the dephasing term. Recall that |±$|\pm\rangle_{j}=\big{(}|0\rangle_{j}\pm|1\rangle_{j}\big{)}/{\sqrt{2}}$ and hence we have $\hat{\sigma}_{j}^{z}|\pm\rangle_{j}=|\mp\rangle_{j}$. The projection of $\mathcal{D}^{\prime}$ to the $l$th layer, $\mathcal{D}_{l}^{\prime}$, couples only $|\lambda_{(l,s)}^{(0)}\rangle\!\rangle$ and $|\lambda_{(l,s\pm 1)}^{(0)}\rangle\!\rangle$, hence it can be written as
where $d_{(l,s)}$ represents the coupling coefficient between the $s$-th and $(s+1)$-th degenerate eigenstates in the $l$-momentum layer. Using the translation symmetry of $\sum_{i}\hat{\sigma}^{z}_{i}$, we have
$d_{(l,s)}$ $=\sum_{j}^{N}\Big{[}\langle E_{s}|\hat{\sigma}^{z}_{j}|E_{s+1}\rangle\langle E_{s+|l|}|\hat{\sigma}^{z}_{j}|E_{s+|l|+1}\rangle\Big{]}=N\Big{[}\langle E_{s}|\hat{\sigma}^{z}_{1}|E_{s+1}\rangle\langle E_{s+|l|}|\hat{\sigma}^{z}_{1}|E_{s+|l|+1}\rangle\Big{]}.$ (S12)
Furthermore, one can show that
$\langle E_{s}|\hat{\sigma}^{z}_{j}|E_{s+1}\rangle=\frac{\binom{N-1}{s}}{\sqrt{\binom{N}{s}\binom{N}{s+1}}}=\frac{1}{N}\sqrt{(N-s)(s+1)}.$ (S13)
Therefore, we have
@@
Next, we consider the effect of self-energy. We assume that the thermal subspace is much larger than the scar subspace, and the energy (initially in scar subspace) that leaks into the thermal subspace cannot return. Thus, we have that the off-diagonal term, $\langle E_{s}|\hat{\Sigma}|E_{s^{\prime}}\rangle$ with $s\neq s^{\prime}$, is much smaller than the diagonal term, $\langle E_{s}|\hat{\Sigma}|E_{s}\rangle$. $\langle E_{s}|\hat{\Sigma}|E_{s}\rangle$ is calculated by Eq. (S7)-(S8). Here, we use approximation as $\langle E_{s}|\hat{\Sigma}|E_{s}\rangle\approx\langle E_{s}|\hat{\Sigma}(E=(-N+2s)J)|E_{s}\rangle$ and obtained the results by solving the self-consistence equation. The latter diagonal element can be written as
@@
$\mathcal{E}_{l}$ $=i\times\mathrm{Diag}\Big{[}\Sigma_{(l,0)},\ \Sigma_{(l,1)},\ \cdots,\Sigma_{(l,N-|l|)}\Big{]},\quad\Sigma_{(l,s)}=\langle E_{s}|\hat{\Sigma}|E_{s}\rangle-\langle E_{s+l}|\hat{\Sigma}|E_{s+l}\rangle.$ (S15)
@@
Finally, we obtain the first-order correction to the eigenvalues $\lambda_{l}^{(1)}$ by diagonalizing the matrix $\big{(}\gamma_{\mathrm{eff}}\mathcal{D}_{l}^{\prime}+\mathcal{E}_{l}\big{)}$ defined in Eq. (S11) and (S15). As an example, consider the SCL model with $N=5$ and the layers $l=l^{\prime\prime}$ with $l^{\prime\prime}=-N,-N+2,\cdots,N$. For example, layer $l=3$ holds triple degeneracy, while layer $l=1$ holds five-fold degeneracy. The corresponding first-order eigenvalues are given by:
@@
  • For the exact scar case $\mathcal{E}=0$, we have $\lambda^{(1)}_{l=3}=0,\pm\frac{4\sqrt{5}}{5}\gamma_{\mathrm{eff}}$ and $\lambda^{(1)}_{l=1}=0,\pm\frac{2\sqrt{10}}{5}\gamma_{\mathrm{eff}},\pm\frac{2\sqrt{46}}{5}\gamma_{\mathrm{eff}}$. The Liouvillean eigenvalues shift along the real axis whenever the dissipative factor $\gamma_{\mathrm{eff}}$ is non-zero;
@@
  • For the approximate scar case $\mathcal{E}\neq 0$, the superoperator of the perturbation is more complicated and in the layers $l=l^{\prime\prime}$ takes the form
@@
    Here we use $\langle E_{s}|\hat{\Sigma}|E_{s}\rangle=-\langle E_{N-s}|\hat{\Sigma}|E_{N-s}\rangle$ because the SCL model holds inversion symmetry. Analogous expressions for the layers $l=l^{\prime}$ are:

### v3

@@ Abstract
$\hat{H}=\hat{S}_{0}\oplus\hat{T}+\hat{V}$, (S1)
where $\hat{S}_{0}$ and $\hat{T}$ denote the Hamiltonian of scar and thermal subsystems, respectively, while $\hat{V}$ represents the coupling between them. This general Hamiltonian includes a large class of QMBS models [11].
The effect of the thermal subspace $\hat{T}$ can be incorporated by defining the effective Hamiltonian for the scar subspace as $\hat{S}=\hat{S}_{0}+\hat{\Sigma}$, where $\hat{\Sigma}$ is the self-energy from the thermal subspace. If we neglect the coupling between the bath and the thermal subsystem, the resulting Liouvillian can be written as
@@
Here, the Liouvillian of the exact scar subspace is denoted by $\mathcal{S}_{0}$, while the self-energy and the dissipator are
@@
The self-energy can be obtained via the non-equilibrium Green’s function as shown below. When $\mathcal{E}$ is zero, the problem reduces to the exact scar case. The effective dephasing strength, $\gamma_{\mathrm{eff}}$, in Eq. (S2) is model-dependent.
The dephasing-free Liouvillian of the exact scar subspace is given by
$\mathcal{S}_{0}={-iJ\Big(\hat{S}_{0}\otimes\hat{I}-\hat{I}\otimes\hat{S}_{0}\Big)},$ (S4)
@@
$|\lambda_{(l,s)}^{(0)}\rangle\!\rangle=|E_{l+s}^{\mathrm{S}}\rangle\otimes|E_{s}^{\mathrm{S}}\rangle,\quad\quad\quad\quad\lambda_{l}^{(0)}=i(E_{l+s}^{\mathrm{S}}-E_{s}^{\mathrm{S}})=2ilJ.$ (S5)
The eigenmodes are labeled using indices $l=-N,-N+1,\cdots,N$ and $s=0,1,2,\cdots,N-|l|$. $E_{s}^{\mathrm{S}}$ and $|E_{s}^{\mathrm{S}}\rangle$ is the scar eigenvalue and eigensate for the Hamitonian of $\hat{S}_{0}$, respectively.
Figure S1: (a) Schematic of a general open quantum many-body system with approximate scars. The scar and thermal subsystems lose their quantum coherence to the bath with the dephasing rate $\gamma$, while interacting via the coupling strength $V$. (b) The Liouvillian spectral transition originates from the competition between self-energy, $\mathcal{E}$, and dissipation $\mathcal{D}$. A set of eigenmodes in Liouville space, denoted by blue concentric circles, are surrounded by the self-energy shell $\mathcal{E}$, due to the thermal subspace. The latter suppresses those dephasing processes (wavy arrows) that do not exceed the boundary of the self-energy shell (yellow circle).
S1.1 A1. The cases with $\mathcal{T}_{-}$ symmetry: CL model
[…]
where $|\pm\rangle_{j}=\big(|{\Uparrow}\rangle_{j}\pm|{\Downarrow}\rangle_{j}\big)/{\sqrt{2}}$ are the spin bases states pointing along the $x$-direction, and $n_{S}=0,1,\ldots,N$ is the number of $|+\rangle$ appearing in the tensor product. In Eq. (S6), we have written the eigenstates belonging to the maximal spin representation.
To explain the structure of Liouvillian QMBS spectra, we set up a degenerate perturbation theory framework. We consider the Liouvillian of the exact scar subspace, Eqs. (S4) and (S6), with eigenvalues
$\lambda_{l}^{(0)}=i(E_{l+s}^{\mathrm{S}}-E_{s}^{\mathrm{S}})=2ilJ,$ (S7)
where $l=-N,-N+1,\cdots,N-1,N$. The eigenvalues $\lambda_{l}$ are equally spaced along the imaginary axis, with the real part equal to zero.
In the regime $\|\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}+\mathcal{E}\|\ll\|\mathcal{S}_{0}\|$, we can use the Liouvillian perturbation theory, in which the traceless Liouvillian eigenvalues for the $l$th layer are approximately given by $\lambda^{\prime}_{l}\approx\lambda^{(0)}_{l}+\lambda^{(1)}_{l}$. Here, the zeroth-order eigenvalues are given by the eigenvalues of the unperturbed system in Eq. (S7), while the first-order corrections to the eigenvalues, $\lambda_{l}^{(1)}$, are obtained by diagonalizing the projection of $\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}+\mathcal{E}$ to the degenerate subspace of the $l$th layer, labeled by states $s=0,1,2,\cdots,N-|l|$.
First, we consider the dephasing term. Recall that $|\pm\rangle_{j}=\big(|0\rangle_{j}\pm|1\rangle_{j}\big)/{\sqrt{2}}$ and hence we have $\hat{\mathbb{S}}_{j}^{z}|\pm\rangle_{j}=|\mp\rangle_{j}$. The projection of $\mathcal{D}^{\prime}$ to the $l$th layer, $\mathcal{D}_{l}^{\prime}$, couples only $|\lambda_{(l,s)}^{(0)}\rangle\!\rangle$ and $|\lambda_{(l,s\pm 1)}^{(0)}\rangle\!\rangle$, hence it can be written as
where $d_{(l,s)}$ represents the coupling coefficient between the $s$-th and $(s+1)$-th degenerate eigenstates in the $l$-momentum layer. Using the translation symmetry of $\sum_{j}\hat{\mathbb{S}}^{z}_{j}$, we have
$d_{(l,s)}$ $=\sum_{j=1}^{N}\Big[\langle E_{s}^{\mathrm{S}}|\hat{\mathbb{S}}^{z}_{j}|E_{s+1}^{\mathrm{S}}\rangle\langle E_{s+|l|}^{\mathrm{S}}|\hat{\mathbb{S}}^{z}_{j}|E_{s+|l|+1}^{\mathrm{S}}\rangle\Big]=N\Big[\langle E_{s}^{\mathrm{S}}|\hat{\mathbb{S}}^{z}_{1}|E_{s+1}^{\mathrm{S}}\rangle\langle E_{s+|l|}^{\mathrm{S}}|\hat{\mathbb{S}}^{z}_{1}|E_{s+|l|+1}^{\mathrm{S}}\rangle\Big].$ (S9)
Furthermore, one can show that
$\langle E_{s}^{\mathrm{S}}|\hat{\mathbb{S}}^{z}_{1}|E_{s+1}^{\mathrm{S}}\rangle=\frac{\binom{N-1}{s}}{\sqrt{\binom{N}{s}\binom{N}{s+1}}}=\frac{1}{N}\sqrt{(N-s)(s+1)}.$ (S10)
Therefore, we have
Next, we consider the effect of self-energy. The Hamiltonian in Eq. (S1) is generally non-integrable and impossible to solve exactly. Hence we employ the non-equilibrium Green’s function approach to model the effect of thermal bulk on the scar subspace, see Ref. [32] for details. This will allow us to obtain analytical insights into the interplay of scarring with the dephasing dissipation. The Green’s function of the scar subspace is expressed as
@@
where $\hat{\Sigma}(E)=\sum_{p_{\mathrm{S}}}\hat{\sigma}_{p_{\mathrm{S}}}(E)$ represents the self-energy matrix, with $\hat{\sigma}_{p_{\mathrm{S}}}(E)=\hat{\Gamma}_{p_{\mathrm{S}}}^{\dagger}\hat{G}_{\mathrm{S}}(E)\hat{\Gamma}_{p_{\mathrm{S}}}\text{ and }\hat{\Gamma}_{p_{\mathrm{S}}}=\sqrt{\gamma_{p_{\mathrm{S}}}}\left|p_{\mathrm{S}}\right\rangle\left\langle p_{\mathrm{S}}\right|$. Here, $|p_{\mathrm{S}}\rangle$ denote the product states in the scar subspace, and $\gamma_{p_{\mathrm{S}}}$ represents the interaction strength between the $|p_{\mathrm{S}}\rangle$ and the thermal subspace. For the QMBS model we consider here, Eq. (S1), $\hat{S}_{0}$ corresponds to a hypercube and $|p_{\mathrm{S}}\rangle$ represent some of the computational basis states, while the remaining basis states $|p_{\mathrm{T}}\rangle$ span the thermal subspace. For example, in the CL model, the vertices of the hypercube are product states of dimers, $\{|p_{\mathrm{S}}\rangle\}=\{\otimes_{j=1}^{N=L/2}|\mathbb{D}_{j}\rangle\}$, where each $|\mathbb{D}_{j}\rangle$ can be either $|\begin{subarray}{c}\downarrow\\ \uparrow\end{subarray}\rangle_{j}$ or $|\begin{subarray}{c}\uparrow\\ \downarrow\end{subarray}\rangle_{j}$. The thermal states $|p_{\mathrm{T}}\rangle$ then involve the remaining basis states that contain $|\begin{subarray}{c}\uparrow\\ \uparrow\end{subarray}\rangle_{j}$ and $|\begin{subarray}{c}\downarrow\\ \downarrow\end{subarray}\rangle_{j}$ states of dimers.
@@
For the CL model, the interaction strength $\gamma_{p_{\mathrm{S}}}$ is proportional to the hopping matrix element $h_{p_{\mathrm{S}}}=\sum_{p_{\mathrm{T}}}\left\langle p_{\mathrm{T}}\right|\hat{V}\left|p_{\mathrm{S}}\right\rangle$ between the scar $|p_{\mathrm{S}}\rangle$ and the thermal subspace $|p_{\mathrm{T}}\rangle$ [32]. The expression for $\hat{\Gamma}$ can be derived analytically:
$\hat{\Gamma}_{\mathrm{}}=\sum_{p_{\mathrm{S}}}\sqrt{\gamma_{p_{\mathrm{S}}}}|p_{\mathrm{S}}\rangle\langle p_{\mathrm{S}}|=\sum_{p_{\mathrm{S}}}|p_{\mathrm{S}}\rangle\langle p_{\mathrm{S}}|\sqrt{\sum_{p_{\mathrm{T}}}\left\langle p_{\mathrm{T}}\right|\hat{V}\left|p_{\mathrm{S}}\right\rangle}=\sum_{p_{\mathrm{S}}}|p_{\mathrm{S}}\rangle\langle p_{\mathrm{S}}|\sqrt{\sum_{j=1}^{N-1}J_{\mathrm{x},j}\hat{P}^{j,j+1}_{p_{\mathrm{S}}}},$ (S14)
@@
where $\hat{P}_{p_{\mathrm{S}}}^{j,j+1}=\langle\mathbb{D}_{j}|\mathbb{D}_{j+1}\rangle$ represents the overlap between two neighboring dimers in $|p_{\mathrm{S}}\rangle$. In the following, we use this expression for $\hat{\Gamma}$ to evaluate the self-energy from the Dyson equation.
@@
We assume that the thermal subspace is much larger than the scar subspace, and the energy (initially in scar subspace) that leaks into the thermal subspace cannot return. Thus, we have that the off-diagonal term, $\langle E_{s}^{\mathrm{S}}|\hat{\Sigma}|E_{s^{\prime}}^{\mathrm{S}}\rangle$ with $s\neq s^{\prime}$, is much smaller than the diagonal term, $\langle E_{s}^{\mathrm{S}}|\hat{\Sigma}|E_{s}^{\mathrm{S}}\rangle$. $\langle E_{s}^{\mathrm{S}}|\hat{\Sigma}|E_{s}^{\mathrm{S}}\rangle$ is calculated by Eq. (S12)-(S13). Here, we use approximation as $\langle E_{s}^{\mathrm{S}}|\hat{\Sigma}|E_{s}^{\mathrm{S}}\rangle\approx\langle E_{s}^{\mathrm{S}}|\hat{\Sigma}(E=(-N+2s)J)|E_{s}^{\mathrm{S}}\rangle$ and obtained the results by solving the self-consistence equation. The latter diagonal element can be written as
@@
$\mathcal{E}_{l}$ $=i\times\mathrm{Diag}\Big[\Sigma_{(l,0)},\ \Sigma_{(l,1)},\ \cdots,\Sigma_{(l,N-|l|)}\Big],\quad\Sigma_{(l,s)}=\langle E_{s}^{\mathrm{S}}|\hat{\Sigma}|E_{s}^{\mathrm{S}}\rangle-\langle E_{s+l}^{\mathrm{S}}|\hat{\Sigma}|E_{s+l}^{\mathrm{S}}\rangle.$ (S15)
@@
Finally, we obtain the first-order correction to the eigenvalues $\lambda_{l}^{(1)}$ by diagonalizing the matrix $\big(\gamma_{\mathrm{eff}}\mathcal{D}_{l}^{\prime}+\mathcal{E}_{l}\big)$ defined in Eq. (S8) and (S15). As an example, consider the CL model with $N=5$ and the layers $l=-N,-N+2,\cdots,N$. For example, layer $l=3$ holds triple degeneracy, while layer $l=1$ holds five-fold degeneracy. The corresponding first-order eigenvalues are given by:
@@
  • For the exact scar case $\mathcal{E}=0$, we have $\lambda^{(1)}_{l=3}=0,\pm\frac{4\sqrt{5}}{5}\gamma_{\mathrm{eff}}$ and $\lambda^{(1)}_{l=1}=0,\pm\frac{2\sqrt{10}}{5}\gamma_{\mathrm{eff}},\pm\frac{2\sqrt{46}}{5}\gamma_{\mathrm{eff}}$. The Liouvillian eigenvalues shift along the real axis whenever the dissipative factor $\gamma_{\mathrm{eff}}$ is non-zero;
@@
  • For the approximate scar case $\mathcal{E}\neq 0$, the superoperator of the perturbation is more complicated and in the layers $l=-N,-N+2,\cdots,N$ takes the form
@@
    Here we use $\langle E_{s}^{\mathrm{S}}|\hat{\Sigma}|E_{s}^{\mathrm{S}}\rangle=-\langle E_{N-s}^{\mathrm{S}}|\hat{\Sigma}|E_{N-s}^{\mathrm{S}}\rangle$ because the CL model holds inversion symmetry. Analogous expressions for the layers $l=-N+1,-N+3,\cdots,N-1$ are: